The sum of the third and twice the first of three consecutive odd integers is 49
Answer by rossiv53(27)

Let x = 1st number
(x+2) = 2nd number
(x+4) = 3rd number
Since the question asks for the sum of twice the first number and the third number, we will add those two values.
(x+4)+2x=49
3x+4=49
3x+4-4=49-4
3x=45
x=15
So 2x = 30. 30 + your 3rd number(19) equals 49.
Good luck!